# Additive approximation algorithm for geodesic centers in $\delta$-hyperbolic graphs

### v1

@@ 1 Introduction
Our algorithm has mainly two stages. In the first stage, we solve the “rooted” version of $(2k-1)$-Geodesic Center, where we require that all isometric paths in the solution has a common end-vertex. Then to reduce the number of isometric paths, we use the shallow pairing property of $\delta$-hyperbolic graphs. See Definition 3.2. Intuitively, this property ensures that the $2k$-many end-vertices of the $12k-$ 1 isometric paths obtained in the first stage can be “paired” to obtain $k$ many isometric paths which together provide an additive $O(\delta)$-approximation algorithm for $k$-Geodesic Center. To the best of our knowledge, the shallow pairing property of $\delta$-hyperbolic graphs was previously unknown and could be of independent interest.
@@ 2 Preliminaries
[…]
@@ 3 Pairings and shallow pairings
In this section we first recall the definition of the pairing property following the terminology of [9]. Then, we introduce a coarse version of this property, called shallow pairing property, and show that this relaxed property holds for $\delta$-hyperbolic graphs.
@@
In general, $\delta$-hyperbolic graphs do not satisfy the pairing property but, as shown below, they satisfy some coarse variant of the pairing property. Before defining this variant, let first define the notion of $\gamma$-shallow pairing
@@ 4.1 An algorithm for the Rooted $k$-Geodesic Center problem
1 $X=V(G)$; $i=0$; while  $i\leq 2k-1$ and $X\neq\emptyset$ do
2       Let $v_{i}\in X$ be a vertex with $d(r,v_{i})\geq d(r,z)$ for all $z\in X$; Let $\sigma_{i}$ be any $(r,v_{i})$-isometric path; $X_{i}=\left\{u\in X:\ \exists\ \text{an $r$-path}\ P\ \text{such that}\ d(u,P)\leq R\ \text{and}\ d(v_{i},P)\leq R\right\}$ $X=X\setminus X_{i}$; $\mathcal{P}=\mathcal{P}\cup\{v_{i}\}$; $\mathcal{C}=\mathcal{C}\cup\{\sigma_{i}\}$. $i=i+1$;
3if $X=\emptyset$ then
@@ 1 $X=V(G)$; $i=0$; while  $i\leq 2k-1$ and $X\neq\emptyset$ do
1 for $v\in V(G)$ do
       Let $R_{v}$ be the smallest $R$ integer for which Algorithm 1 returns a $(v,R+2\delta)$-cover $\mathcal{C}_{v}$ of size $12k-$ 1 . /* The above steps can be implemented by using Algorithm 1 in combination with a binary search on $R_{v}\in\{0,1,\ldots,|V(G)|\}$. */
2
Let $u\in V(G)$ such that $R_{u}=\min\{R_{v}\colon v\in V(G)\}$. return $(\mathcal{C}_{u},R_{u})$.
@@ 4.3 From rooted to non rooted collection of paths
Let $\mathcal{C}_{u}$ be the $(u,R_{u})$-cover of $G$ returned by Algorithm 2 with $G$ and $k$ as input. Let $\pi$ be the even profile consisting of $u$ and every other end vertex of the $u$-paths in $\mathcal{C}_{u}$. Computes a $(2\tau(G)+\frac{1}{2})$-shallow pairing $P$ of $\pi.$ return $\mathcal{S}(P)$.

### v2

@@ 1 Introduction
Our algorithm has mainly two stages. In the first stage, we solve the “rooted” version of $(2k-1)$-Geodesic Center, where we require that all isometric paths in the solution has a common end-vertex. Then to reduce the number of isometric paths, we use the shallow pairing property of $\delta$-hyperbolic graphs. See Definition 3.2. Intuitively, this property ensures that the $2k$-many end-vertices of the $12k-$ 1 isometric paths obtained in the first stage can be “paired” to obtain $k$ many isometric paths which together provide an additive $O(\delta)$-approximation algorithm for $k$-Geodesic Center. We think that the shallow pairing property could also be interesting in itself and for other algorithmic applications.
@@ 2 Preliminaries
$\alpha_{y}$$\leq\delta$$\leq\delta$$m^{\prime}$$\varphi$$m_{x}$$m_{z}$$m_{y}$$x$$z$$x^{\prime}$$z^{\prime}$$y$$y^{\prime}$$\leq\delta$$\alpha_{x}$$\alpha_{z}$
@@ 3 Pairings and shallow pairings
In this section we first recall the definition of the pairing property following the terminology of [9]. Then, we introduce a coarse version of this property, called shallow pairing property [^2] The notion of approximate (shallow) pairing in hyperbolic graphs was defined by Victor Chepoi, who also asked the question about their existence., and show that this relaxed property holds for $\delta$-hyperbolic graphs.
@@
In general, $\delta$-hyperbolic graphs do not satisfy the pairing property but, as shown below, they satisfy some coarse variant of the pairing property. Before defining this variant, let us first define the notion of $\gamma$-shallow pairing
@@ 4.1 An algorithm for the Rooted $k$-Geodesic Center problem
1
[…]
3while  $i\leq 2k-1$ and $X\neq\emptyset$ do
4       Let $v_{i}\in X$ be a vertex with $d(r,v_{i})\geq d(r,z)$ for all $z\in X$;
5      Let $\sigma_{i}$ be any $(r,v_{i})$-isometric path;
6      $X_{i}=\left\{u\in X:\ \exists\ \text{an $r$-path}\ P\ \text{such that}\ d(u,P)\leq R\ \text{and}\ d(v_{i},P)\leq R\right\}$
7      $X=X\setminus X_{i}$;
8      $\mathcal{P}=\mathcal{P}\cup\{v_{i}\}$;
9      $\mathcal{C}=\mathcal{C}\cup\{\sigma_{i}\}$.
10      $i=i+1$;
11
12if $X=\emptyset$ then
@@ 10      $i=i+1$;
1
2for $v\in V(G)$ do
3
4      Let $R_{v}$ be the smallest $R$ integer for which Algorithm 1 returns a $(v,R+2\delta)$-cover $\mathcal{C}_{v}$ of size $12k-$ 1 .
      /* The above steps can be implemented by using Algorithm 1 in combination with a binary search on $R_{v}\in\{0,1,\ldots,|V(G)|\}$. */
5
6Let $u\in V(G)$ such that $R_{u}=\min\{R_{v}\colon v\in V(G)\}$.
7return $(\mathcal{C}_{u},R_{u})$.
8
@@ 4.3 From rooted to non rooted collection of paths
1
2Let $\mathcal{C}_{u}$ be the $(u,R_{u})$-cover of $G$ returned by Algorithm 2 with $G$ and $k$ as input.
3Let $\pi$ be the even profile consisting of $u$ and every other end vertex of the $u$-paths in $\mathcal{C}_{u}$.
4Computes a $(2\tau(G)+\frac{1}{2})$-shallow pairing $P$ of $\pi.$
5return $\mathcal{S}(P)$.
6